### v1

@@ 3.4 IR-maser correlation
For the G323 burst, there is also a clear correlation between the maser flux and the -band flux (cf. Fig. 4 ). Note that although NIR -band photons do not pump the masers, the long-term variation of the integrated maser flux closely follows the light curve, simply because the NIR and MIR maser-exciting radiation vary in the same fashion. A log-log regression yields a power-law exponent of . Thus, the G323 burst strengthens the evidence that Class II methanol maser flares are a reliable tracer of episodic accretion variability of young massive stars.
@@ 3.5 HAWC+ FIR photometry
As stated above, the main objective of the SOFIA/HAWC+ observations is to constrain the burst energy by assessing the strength of a possible FIR afterglow due to the burst. This was done by comparing the pre-burst fluxes from the HPPSC with those measured with HAWC+.
Although the field sizes of the 53, 62, and 89 m bands do not cover additional HPPSC sources, the larger fields of the 154 and 214 m bands comprise one or more. The detection of the source HPPSC160A_J152920.2-562522 in the 154 m image provides the opportunity to check the flux calibration. PSF photometry of this object revealed that its HAWC+ flux exceeds the PACS flux by (157)%. Consequently, this was taken into account in the analysis. Furthermore, the HAWC+ 62 and 154 m fluxes were interpolated to the PACS wavelengths using a polynomial approximation of the post-burst FIR SED established from all HAWC+ bands.
@@
The HAWC+ photometry is given in Table 7 along with beam sizes, as well as image and deconvolved FWHMs. The FWHMs of the image profile were derived by fitting an ellipse to the cut at 50% peak level. The uncertainty of the length of the axis ranges from 05 to 15, while that of the position angles is 13°. The deconvolved sizes are based on the geometric mean of the major and minor axes, from which the beam size was subtracted in quadrature.
The post-burst SED is plotted in Fig. 9 in orange (HAWC+ filters), with the interpolation to the PACS wavelengths in red. The pre-burst SED is shown in blue for comparison. The inset shows a zoom-in on the region of interest.
@@ 3.6 AO and thermal IR imagery
The pre-burst NACO -band image is shown in Fig. 10. A logarithmic scale is used to display the lower surface brightness. The black diamond marks the peak position of 1.4 mm emission, based on the ALMAGAL data obtained with the ALMA 12-m array. This probably corresponds to the location of the MYSO. As mentioned in the caption of Fig. 5, the contours of the emission of the band are superimposed in red. The two scattering blobs are located south-east of the -band peak. The western one coincides with the position of the GAIA source. The white contours delineate the 19 GHz radio continuum (Murphy et al., 2010), which is centered on the MYSO. The remaining symbols, filled circles with error bars, mark the positions of 6.7 GHz masers (blue - Caswell, 1997, red - Green et al., 2015), as well as the width and orientation of the ISAAC slit (yellow). The tightly winded circular -band contours at the bottom are due to a foreground star that is barely visible at their center.
@@ 3.7 K-band spectroscopy
Table 1 lists line fluxes, full width at half maximum (), peak radial velocities () and full width at zero intensity () of Br and H2 along the three different regions extracted from the spectral image, namely at the source, toward the NE and SW of the source. The H2 emission, only detected in the NE region, is spectrally unresolved (2.4 Å or 34 km s-1) and its maximum radial velocity is around 0 km s-1 or slightly red-shifted (44 km s-1). The H2 kinematics points to fluorescent emission from the HCHii region around the source. On the other hand, the kinematics of the Br line is more complex. The line is spectrally resolved in the three extracted regions. It has a radial velocity close to 0 km s-1 on source, whereas the SW region is slightly blueshifted (-145 km s-1) whereas the NE region has three different velocity components at -28, 2, and 32 km s-1. Furthermore, the values of the line in the three regions range from 150 to 170 km s-1. Spectroscopy was performed during the light-curve rise, close to the NIR peak of the outburst. At that time, no P-Cygni profile was present in the Br line. This feature is commonly attributed to an emerging ionized wind. Its absence in the G323 spectrum may indicate that the launching of the wind is not a prompt process, but delayed with respect to the temporal evolution of the burst. This was observed for the S255IR-NIRS3 outburst (Cesaroni et al., 2018).
@@ 3.8 The ALMA view of G323
The contours of the intensity map of the blue- and redshifted 13CO(2–1) emission are shown in Fig. 14, superimposed on a image taken in mid-2015 along with the 1.4 mm continuum emission. The latter appears as a small blob in the white area of the saturated central region. The synthesized beam size () of the 13CO(2–1) observations is shown on the bottom left. The line emission was integrated in the LSR velocity ranges from 7 to 18 km for the redshifted gas and from -7 to -15 km for the blueshifted gas. The contours begin at 5 and increase in steps of 2 where is 0.18 and 0.13 for the redshifted and blueshifted lobes, respectively.
Remarkably, the blueshifted gas appears to coincide with the dusty ridge west of the MYSO while the redshifted emission toward the east is co-spatial with the NIR nebulosity. This is not typical for outflow cavities, which are usually devoid of dust and scatter from their walls. If the CO emission traces an outflow of a source seen almost pole-on, its bipolar lobes should overlap more strongly than is the case for G323. Moreover, considerable bending seems to be present. Although this challenges the outflow interpretation and possibly argues for bulk-gas motion, a deflection of the blueshifted lobe at the western dust rim cannot be ruled out. Unfortunately, the ALMA 1.4 mm map is not sufficiently sensitive to shed light on the dust distribution at larger scales.
Inspection of the SPITZER/IRAC images does not provide evidence for a bipolar outflow as well. Unlike other MYSOs, for which outflows were found (e.g., Caratti o Garatti et al., 2015), the IRAC2 image, which covers a wealth of shock-excited H2 lines (e.g., Noriega-Crespo et al., 2004), does not show typical bow-like emission features on either side of the source.
@@ 4.1 The static pre-burst model
We used MRN dust (Mathis et al., 1977), which consists of compact, homogeneous, and spherical grains with a composition of 62.5% silicate and 37.5% graphite, where we took the optical properties from Weingartner & Draine (2001). The size distribution can be described by a power law with with grain sizes ranging from to .
All parameters were sampled from the ranges given in Table 9. The protostellar luminosity includes all possible mechanisms of energy release (fusion, contraction, accretion). Throughout the paper the term ‘luminosity’ represents the bolometric one, obtained by integrating the SED over the whole wavelength range observed. The inner radius of the circumstellar disk is governed by the dust sublimation radius (assuming ).
For our models, we cut the envelope at an outer radius , which was sampled between (). This is in the range of the size of a typical cloud core. The extent of the FWHM in the ALMA images is smaller by a factor of 2 to 10. Due to the high resolution, the interferometer may not be sensitive to the extended component of the emission. However, most of the emission should be covered by ALMA. The extent of the prompt LE is greater than the outer radius of our models by a factor of . However, the density in the outermost parts of the parent cloud core is probably pretty low.
@@
The G323 pre-burst SED is well covered from NIR to (sub)mm. The number of free parameters is as large as 11 and there are degeneracies among the models. Therefore, we do not use just the best model, but rely on the 10 best models instead. This approach was applied already in Stecklum et al. (2021). All parameters of the ten best pre-burst models (blue) are averaged using weights according to the respective value. For linear sampled parameters, we use the arithmetic mean, whereas for logarithmically sampled parameters, it is the geometric mean. The resulting mean model, i.e., the model where all parameters are mean values, is colored red. The mean values can be found in the right column of Table 9. The table with the 10 best models is given in the Appendix (Table 11). Note that the best models point to an almost face-on view. The inclination is smaller than the opening angle of the cavity, which implies that the optical depth toward the object is minimized.
In addition to the mean model, two additional settings are included (Tmin and Tmax). These settings minimize/maximize the afterglow timescales and compromise parameters, which are in agreement with the pre-burst fit within the given confidence intervals. The Tmin setting has a lower dust content, and a wider opening angle of the cavity compared to the mean setting. The Tmax setting has a higher dust content, and a smaller cavity opening angle. The parameters of all three models are summarized in Table 10, where only those that differ are given. We changed only those parameters that are considered to strongly affect the duration of the afterglow. All other parameters are the same as for the mean model (right column of Table 9). Constant parameters are the protostellar luminosity, inclination, and most of the disk parameters, since their impact is small or unclear. The Tmin, and Tmax settings are meant to give limits on the burst energy that take into account that the afterglow depends not only on the burst energy but also on the local dust distribution. Both the mean, and the Tmax models agree well with the pre-burst SED (their is comparable or better than that of the best-fit model). The Tmin model has too little cold dust and cannot reproduce the FIR fluxes. However, TDRT models were also performed for this model.
@@ 4.2.1 TORUS-Code
To model the temporal evolution of the SED, the TORUS radiative transfer code was used in its time-dependent mode. The time-dependent algorithm follows the random walk of photon packets as they propagate over a time step, with the energy absorption rates for each grid cell calculated using a MC estimator and the energy emission rate calculated from the local dust temperature. The TDRT method is described in detail in Harries (2011) and has been benchmarked against several standard thermodynamic test problems. Unlike traditional methods, such as flux-limited diffusion, the code is able to simultaneously and accurately treat both the free-streaming and diffusive regimes. The role of the time step is crucial to the accuracy of the calculation. It needs to be short enough for the luminosity variation to remain tractable, but long enough to avoid creating too much computational overhead (Harries et al., 2019). We used a time step of , which was verified with an interval shorter by a factor 7. With this choice, the envelope heating, which is responsible for most of the FIR radiation, is traced very well. But for the disk, a shorter time step is better. TORUS is a very versatile code with several microphysics modules from which we only apply those to calculate the dust radiative equilibrium (static RT) and non-equilibrium (TDRT). The former was used, for example by Esau et al. (2014) while the latter was validated against the Pascucci disk benchmark (Pascucci et al., 2004).
@@ 4.2.2 Assumptions and constraints
First, we examine whether the energy radiated away by dust is a good proxy for the total energy released by the burst. The two main coolants for YSOs are molecular lines, in particular from CO, and dust radiation. Dust grains are solids, i.e., thermal emitters. Their opacities cover more than four orders of magnitude in wavelength of the electromagnetic spectrum (e.g., Min, 2015) and are thus much more capable of affecting radiative heat transfer than gas opacities. In fact, the luminosity of the FIR molecular cooling lines of MYSOs is only a tiny fraction (, Karska et al., 2014) of the dust luminosity. The same holds for the energy release by free-free emission from a compact Hii region (R. Cesaroni, priv. comm.). Therefore, the luminosity estimate derived from the SED can be considered as the total value.
@@
For TDRT modeling, we use the most simple approach possible, namely, we include only heating and cooling but omit any kind of density variation as well as any changes in the disk chemistry. The latter is generally not met. Instead, the burst will most likely sublimate the innermost disk, change the chemistry of the disk, and shift the ice line outward (e.g. Lee, 2007, Visser & Bergin, 2012, Vorobyov et al., 2013, Rab et al., 2017).
In principle, dust sublimation is implemented in TORUS but it requires a time step of less than 1 d to remove the disk on typical disk dispersal timescales. However, the impact on the evolution of the (F)IR luminosity will be minor. Therefore, we use a longer time step, where we move the inner radius outward to ensure that the dust does not become hotter than . We shift it to 60 au, which is according to (Whitney et al., 2004, Eq. 1, with K), where we kept the disk mass constant.
@@ 4.3 The dynamic SED of the mean model
Both and increase with wavelength. This is expected since the radiation at longer wavelengths can be attributed to regions more distant from the star (i.e., colder regions). Interestingly, the delay between NIR and FIR is almost 2 (and 3) years for (and ) at . For comparison, the light travel time from the source to the outer edge of the grid is only . The delay between the peak of the burst profile (1.2 yr after onset) and that in the FIR is by far more than what could be explained by geometrical/projection effects and distinct spatial origins alone. It clearly indicates a measurable slowdown of the energy transfer toward the FIR-emitting regions by numerous absorption and re-emission processes owing to the high optical depths in between the protostar and FIR-emitting regions. Toward longer wavelengths () both curves flatten. This is expected because at some point even the coldest and most distant regions are ‘processed’ and we are reaching the Rayleigh-Jeans tail of the emission from the coldest dust with . Furthermore, these regions are generally optically thin at these wavelengths.
@@ 4.4.1 Burst energy estimate for the mean model
To constrain the burst energy, we performed 7 simulations with burst energies between (0.2 and . The corresponding burst templates (simulation input) are displayed in the left panel of Fig. 17 with the different burst profiles color-coded.
@@
The ratios can be fitted with linear functions in the given range (indicated by the colored solid lines), whereby the y-axis intercept equals 1, as no energy input means no flux increase.
@@ 4.4.2 Limits on the burst energy
We ran 7 simulations for the Tmin setting with burst energies between (2.3 and and 9 simulations for the Tmax setting with burst energies between (0.05 and .
@@
For the Tmax setting, the ratio at m exceeds the one at m. This is in agreement with the dynamic SED (Fig. 19), which shows a strong increase of the indicated timescales in between and m.
Again, the best estimate for the burst energy is indicated by vertical lines. It amounts to for Tmin and for Tmax in units of . For the Tmin and Tmax settings, the values of are about a factor 13 above and 6 below the mean model respectively. We emphasize that the higher value is not likely, as the pre-burst fit of this model is much worse than for the other two settings. However, it may still serve as a conservative upper limit. A summary of the derived burst parameters is provided in Table 4.
@@
In addition to the HAWC+ flux density ratios, the models can be compared with the measurements. Similar to the HAWC+ fit, we used the flux density ratios. Fig. 21 shows the ratio curves for all 3 configurations (from left to right) and all burst energies (color-coded). The observation is shown in red. For the Tmin configuration and the highest burst energies, the flux does not return to its pre-burst level at the end of the burst. This is probably caused by dust that becomes hotter than . We shifted the innermost radius to 60 au () to avoid too high dust temperatures for all models. This is not sufficient for the most energetic bursts. All bursts overestimate the observed ratio (red) for the Tmin setting. This is another indication that the burst energy estimate obtained with this setting is too large. We emphasize that the ratio depends only slightly on the setting. For nearly the same burst energy for Tmin and the mean setting, the best match is reached with the curve, despite the vastly different afterglows. Therefore, the ratio provides a good measure for the burst energy. For both the mean and the Tmax setting, a good agreement is reached for a burst with an energy of . This value lies well within the determined range and would imply a dust configuration in the range between mean and Tmax.
Probably the best estimate of the burst energy is , which is based on the value and covers the entire range spanned by the mean and Tmax settings. This would correspond to a peak value of , and hence an increase of the luminosity of the protostar by a factor (or ).

### v2

@@ 3.4 IR-maser correlation
For the G323 burst, there is also a clear correlation between the maser flux and the -band flux (cf. Fig. 4 ). We note that although NIR -band photons do not pump the masers, the long-term variation of the integrated maser flux closely follows the light curve, simply because the NIR and MIR maser-exciting radiation vary in the same fashion. A log-log regression yields a power-law exponent of . Thus, the G323 burst strengthens the evidence that Class II methanol maser flares are a reliable tracer of episodic accretion variability of young massive stars.
@@ 3.5 HAWC+ FIR photometry
As stated above, the main objective of the SOFIA/HAWC+ observations is to constrain the burst energy by assessing the strength of a possible FIR afterglow due to the burst. This was done by comparing the pre-burst fluxes from the HPPSC with those measured with HAWC+. Although the field sizes of the 53, 62, and 89 m bands do not cover additional HPPSC sources, the larger fields of the 154 and 214 m bands comprise one or more. The detection of the source HPPSC160A_J152920.2-562522 in the 154 m image provides the opportunity to check the flux calibration. PSF photometry of this object revealed that its HAWC+ flux exceeds the PACS flux by (157)%. Consequently, this was taken into account in the analysis. Furthermore, the HAWC+ 62 and 154 m fluxes were interpolated to the PACS wavelengths using a polynomial approximation of the post-burst FIR SED established from all HAWC+ bands.
@@
The HAWC+ photometry is given in Table 7 along with beam sizes, as well as image and deconvolved FWHMs. The FWHMs of the image profile were derived by fitting an ellipse to the cut at 50% peak level. The uncertainty of the length of the axis ranges from 05 to 15, while that of the position angles is 13°. The deconvolved sizes are based on the geometric mean of the major and minor axes, from which the beam size was subtracted in quadrature. The post-burst SED is plotted in Fig. 9 in orange (HAWC+ filters), with the interpolation to the PACS wavelengths in red. The pre-burst SED is shown in blue for comparison. The inset shows a zoom-in on the region of interest.
@@ 3.6 AO and thermal IR imagery
The pre-burst NACO -band image is shown in Fig. 10. A logarithmic scale is used to display the lower surface brightness. The black diamond marks the peak position of 1.4 mm emission, based on the ALMAGAL data obtained with the ALMA 12-m array. This probably corresponds to the location of the MYSO. As mentioned in the caption of Fig. 5, the contours of the emission of the band are superimposed in red. The two scattering blobs are located southeast of the -band peak. The western one coincides with the position of the GAIA source. The white contours delineate the 19 GHz radio continuum (Murphy et al., 2010), which is centered on the MYSO. The remaining symbols, filled circles with error bars, mark the positions of 6.7 GHz masers (blue - Caswell, 1997, red - Green et al., 2015), as well as the width and orientation of the ISAAC slit (yellow). The tightly winded circular -band contours at the bottom are due to a foreground star that is barely visible at their center.
@@ 3.7 K-band spectroscopy
Table 1 lists line fluxes, full width at half maximum (), peak radial velocities () and full width at zero intensity () of Br and H2 along the three different regions extracted from the spectral image, namely at the source, toward the NE and SW of the source. The H2 emission, only detected in the NE region, is spectrally unresolved (2.4 Å or 34 km s-1) and its maximum radial velocity is around 0 km s-1 or slightly redshifted (44 km s-1). The H2 kinematics points to fluorescent emission from the HCHii region around the source. On the other hand, the kinematics of the Br line is more complex. The line is spectrally resolved in the three extracted regions. It has a radial velocity close to 0 km s-1 on source, whereas the SW region is slightly blueshifted (-145 km s-1) whereas the NE region has three different velocity components at -28, 2, and 32 km s-1. Furthermore, the values of the line in the three regions range from 150 to 170 km s-1. Spectroscopy was performed during the light-curve rise, close to the NIR peak of the outburst. At that time, no P-Cygni profile was present in the Br line. This feature is commonly attributed to an emerging ionized wind. Its absence in the G323 spectrum may indicate that the launching of the wind is not a prompt process, but delayed with respect to the temporal evolution of the burst. This was observed for the S255IR-NIRS3 outburst (Cesaroni et al., 2018).
@@ 3.8 The ALMA view of G323
The contours of the intensity map of the blueshifted and redshifted 13CO(2–1) emission are shown in Fig. 14, superimposed on a image taken in mid-2015 along with the 1.4 mm continuum emission. The latter appears as a small blob in the white area of the saturated central region. The synthesized beam size () of the 13CO(2–1) observations is shown on the bottom left. The line emission was integrated in the LSR velocity ranges from 7 to 18 km for the redshifted gas and from -7 to -15 km for the blueshifted gas. The contours begin at 5 and increase in steps of 2 where is 0.18 and 0.13 for the redshifted and blueshifted lobes, respectively. Remarkably, the blueshifted gas appears to coincide with the dusty ridge west of the MYSO while the redshifted emission toward the east is co-spatial with the NIR nebulosity. This is not typical for outflow cavities, which are usually devoid of dust and scatter from their walls. If the CO emission traces an outflow of a source seen almost pole-on, its bipolar lobes should overlap more strongly than is the case for G323. Moreover, considerable bending seems to be present. Although this challenges the outflow interpretation and possibly argues for bulk-gas motion, a deflection of the blueshifted lobe at the western dust rim cannot be ruled out. Unfortunately, the ALMA 1.4 mm map is not sufficiently sensitive to shed light on the dust distribution at larger scales. Inspection of the SPITZER/IRAC images does not provide evidence for a bipolar outflow as well. Unlike other MYSOs, for which outflows were found (e.g., Caratti o Garatti et al., 2015), the IRAC2 image, which covers a wealth of shock-excited H2 lines (e.g., Noriega-Crespo et al., 2004), does not show typical bow-like emission features on either side of the source.
@@ 4.1 The static pre-burst model
We used MRN dust (Mathis et al., 1977), which consists of compact, homogeneous, and spherical grains with a composition of 62.5% silicate and 37.5% graphite, where we took the optical properties from Weingartner & Draine (2001). The size distribution can be described by a power law with with grain sizes ranging from to . All parameters were sampled from the ranges given in Table 9. The protostellar luminosity includes all possible mechanisms of energy release (fusion, contraction, accretion). Throughout the paper the term ‘luminosity’ represents the bolometric one, obtained by integrating the SED over the whole wavelength range observed. The inner radius of the circumstellar disk is governed by the dust sublimation radius (assuming ). For our models, we cut the envelope at an outer radius , which was sampled between (). This is in the range of the size of a typical cloud core. The extent of the FWHM in the ALMA images is smaller by a factor of two to ten. Due to the high resolution, the interferometer may not be sensitive to the extended component of the emission. However, most of the emission should be covered by ALMA. The extent of the prompt LE is greater than the outer radius of our models by a factor of . However, the density in the outermost parts of the parent cloud core is probably pretty low.
@@
The G323 pre-burst SED is well covered from NIR to (sub)mm. The number of free parameters is as large as 11 and there are degeneracies among the models. Therefore, we do not use just the best model, but rely on the ten best models instead. This approach was applied already in Stecklum et al. (2021). All parameters of the ten best pre-burst models (blue) are averaged using weights according to the respective value. For linear sampled parameters, we use the arithmetic mean, whereas for logarithmically sampled parameters, it is the geometric mean. The resulting mean model, that is, the model where all parameters are mean values, is colored red. The mean values can be found in the right column of Table 9. The table with the ten best models is given in the Appendix (Table 11). Note that the best models point to an almost face-on view. The inclination is smaller than the opening angle of the cavity, which implies that the optical depth toward the object is minimized.
In addition to the mean model, two additional settings are included (Tmin and Tmax). These settings minimize and maximize the afterglow timescales and compromise parameters, which are in agreement with the pre-burst fit within the given confidence intervals. The Tmin setting has a lower dust content, and a wider opening angle of the cavity compared to the mean setting. The Tmax setting has a higher dust content, and a smaller cavity opening angle. The parameters of all three models are summarized in Table 10, where only those that differ are given. We changed only those parameters that are considered to strongly affect the duration of the afterglow. All other parameters are the same as for the mean model (right column of Table 9). Constant parameters are the protostellar luminosity, inclination, and most of the disk parameters, since their impact is small or unclear. The Tmin, and Tmax settings are meant to give limits on the burst energy that take into account that the afterglow depends not only on the burst energy but also on the local dust distribution. Both the mean, and the Tmax models agree well with the pre-burst SED (their is comparable or better than that of the best-fit model). The Tmin model has too little cold dust and cannot reproduce the FIR fluxes. However, TDRT models were also performed for this model.
@@ 4.2.1 TORUS-Code
To model the temporal evolution of the SED, the TORUS radiative transfer code was used in its time-dependent mode. The time-dependent algorithm follows the random walk of photon packets as they propagate over a time step, with the energy absorption rates for each grid cell calculated using a MC estimator and the energy emission rate calculated from the local dust temperature. The TDRT method is described in detail in Harries (2011) and has been benchmarked against several standard thermodynamic test problems. Unlike traditional methods, such as flux-limited diffusion, the code is able to simultaneously and accurately treat both the free-streaming and diffusive regimes. The role of the time step is crucial to the accuracy of the calculation. It needs to be short enough for the luminosity variation to remain tractable, but long enough to avoid creating too much computational overhead (Harries et al., 2019). We used a time step of , which was verified with an interval shorter by a factor 7. With this choice, the envelope heating, which is responsible for most of the FIR radiation, is traced very well. But for the disk, a shorter time step is better. TORUS is a very versatile code with several microphysics modules from which we only apply those to calculate the dust radiative equilibrium (static RT) and nonequilibrium (TDRT). The former was used, for example by Esau et al. (2014) while the latter was validated against the Pascucci disk benchmark (Pascucci et al., 2004).
@@ 4.2.2 Assumptions and constraints
First, we examine whether the energy radiated away by dust is a good proxy for the total energy released by the burst. The two main coolants for YSOs are molecular lines, in particular from CO, and dust radiation. Dust grains are solids, which implies that they are thermal emitters. Their opacities cover more than four orders of magnitude in wavelength of the electromagnetic spectrum (e.g., Min, 2015) and are thus much more capable of affecting radiative heat transfer than gas opacities. In fact, the luminosity of the FIR molecular cooling lines of MYSOs is only a tiny fraction (, Karska et al., 2014) of the dust luminosity. The same holds for the energy release by free-free emission from a compact Hii region (R. Cesaroni, priv. comm.). Therefore, the luminosity estimate derived from the SED can be considered as the total value.
@@
For TDRT modeling, we use the most simple approach possible, namely, we include only heating and cooling but omit any kind of density variation as well as any changes in the disk chemistry. The latter is generally not met. Instead, the burst will most likely sublimate the innermost disk, change the chemistry of the disk, and shift the ice line outward (e.g., Lee, 2007, Visser & Bergin, 2012, Vorobyov et al., 2013, Rab et al., 2017). In principle, dust sublimation is implemented in TORUS but it requires a time step of less than 1 d to remove the disk on typical disk dispersal timescales. However, the impact on the evolution of the (F)IR luminosity will be minor. Therefore, we use a longer time step, where we move the inner radius outward to ensure that the dust does not become hotter than . We shift it to 60 au, which is according to (Whitney et al., 2004, Eq. 1, with K), where we kept the disk mass constant.
@@ 4.3 The dynamic SED of the mean model
Both and increase with wavelength. This is expected since the radiation at longer wavelengths can be attributed to regions more distant from the star (i.e., colder regions). Interestingly, the delay between NIR and FIR is almost two (and three) years for (and ) at . For comparison, the light travel time from the source to the outer edge of the grid is only . The delay between the peak of the burst profile (1.2 yr after onset) and that in the FIR is by far more than what could be explained by geometrical/projection effects and distinct spatial origins alone. It clearly indicates a measurable slowdown of the energy transfer toward the FIR-emitting regions by numerous absorption and re-emission processes owing to the high optical depths in between the protostar and FIR-emitting regions. Toward longer wavelengths () both curves flatten. This is expected because at some point even the coldest and most distant regions are ‘processed’ and we are reaching the Rayleigh-Jeans tail of the emission from the coldest dust with . Furthermore, these regions are generally optically thin at these wavelengths.
@@ 4.4.1 Burst energy estimate for the mean model
To constrain the burst energy, we performed seven simulations with burst energies between (0.2 and . The corresponding burst templates (simulation input) are displayed in the left panel of Fig. 17 with the different burst profiles color-coded.
@@
The ratios can be fitted with linear functions in the given range (indicated by the colored solid lines), whereby the y-axis intercept equals one, as no energy input means no flux increase.
@@ 4.4.2 Limits on the burst energy
We ran seven simulations for the Tmin setting with burst energies between (2.3 and and nine simulations for the Tmax setting with burst energies between (0.05 and .
@@
For the Tmax setting, the ratio at m exceeds the one at m. This is in agreement with the dynamic SED (Fig. 19), which shows a strong increase of the indicated timescales in between and m. Again, the best estimate for the burst energy is indicated by vertical lines.
It amounts to for Tmin and for Tmax in units of .
For the Tmin and Tmax settings, the values of are about a factor 13 above and a factor six below the mean model respectively.
We emphasize that the higher value is not likely, as the pre-burst fit of this model is much worse than for the other two settings. However, it may still serve as a conservative upper limit. A summary of the derived burst parameters is provided in Table 4.
@@
In addition to the HAWC+ flux density ratios, the models can be compared with the measurements. Similar to the HAWC+ fit, we used the flux density ratios. Fig. 21 shows the ratio curves for all three configurations (from left to right) and all burst energies (color-coded). The observation is shown in red. For the Tmin configuration and the highest burst energies, the flux does not return to its pre-burst level at the end of the burst. This is probably caused by dust that becomes hotter than . We shifted the innermost radius to 60 au () to avoid too high dust temperatures for all models. This is not sufficient for the most energetic bursts. All bursts overestimate the observed ratio (red) for the Tmin setting. This is another indication that the burst energy estimate obtained with this setting is too large. We emphasize that the ratio depends only slightly on the setting. For nearly the same burst energy for Tmin and the mean setting, the best match is reached with the curve, despite the vastly different afterglows. Therefore, the ratio provides a good measure for the burst energy. For both the mean and the Tmax setting, a good agreement is reached for a burst with an energy of . This value lies well within the determined range and would imply a dust configuration in the range between mean and Tmax. Probably the best estimate of the burst energy is , which is based on the value and covers the entire range spanned by the mean and Tmax settings. This would correspond to a peak value of , and hence an increase of the luminosity of the protostar by a factor (or ).